# Jin Li    Ye Luo    Xiaowei Zhang

### v1

@@
Seesaw Experimentation: A/B Tests with Spillovers thanks: Contact information: Jin Li (jli1@hku.hk), Ye Luo (kurtluo@hku.hk), Faculty of Business and Economics, The University of Hong Kong, Pokfulam Road, Hong Kong SAR. Xiaowei Zhang (xiaoweiz@ust.hk), Department of Industrial Engineering and Decision Analytics, The Hong Kong University of Science and Technology, Clear Water Bay, Hong Kong SAR.
@@
(November 4, 2024)
This paper examines how a firm’s performance can decline despite consistently implementing successful A/B test innovations—a phenomenon we term “seesaw experimentation.” While these innovations improve the measured primary dimension, they create negative externalities in unmeasured secondary dimensions that exceed the gains. Using a multivariate normal distribution model, we identify the conditions for this decline and propose positive hurdle rates as a solution. Our analysis shows how to set optimal hurdle rates to best mitigate these negative externalities and provides practical guidance for experimental design by demonstrating how these rates should vary with underlying parameters.
@@ 1 Introduction
A/B testing has emerged as a fundamental pillar of modern business strategies, with firms leveraging these experiments to drive innovation, enhance user experiences, and boost revenue growth (Thomke (2020), Koning, Hasan and Chatterji (2022), Kohavi, Tang and Xu (2020)). Tech giants like Google and Amazon now deploy thousands of such tests annually, embedding data-driven decision-making into their innovation processes.
At the heart of A/B testing lies a crucial challenge: selecting the right metric for evaluation. This decision is inherently complex because firm performance spans multiple dimensions, and innovations often generate opposing effects across these dimensions. When a firm reduces advertisements to enhance user experience, it may sacrifice revenue. When it boosts online sales, offline purchases might decline.
These spillovers reflect fundamental trade-offs in business performance. [^1] We use spillover, rather than negative spillover, for expositional ease. Managing them poses a significant challenge, particularly for large organizations where the sheer volume of tests makes interdepartmental coordination impractical. Resource constraints further complicate this challenge, as firms often cannot measure all potential effects, especially long-term impacts that resist easy quantification.
This paper examines how spillover effects shape the dynamics of experimentation. These effects arise when innovations improve performance in a primary dimension—typically aligned with current strategic priorities—while generating negative impacts in secondary dimensions. As business environments evolve, so do strategic priorities: today’s focus on revenue might shift to user growth tomorrow. Consequently, even as firms accumulate successful A/B tests in their primary metrics, their overall performance may stagnate or decline.
We call this phenomenon seesaw experimentation—where the cumulative impact of successful A/B tests fails to generate overall business improvement. Like a seesaw, gains in one dimension are counterbalanced by losses in others, creating oscillation between different performance metrics without meaningful progress.
Using a multivariate normal distribution framework, we formally demonstrate how seesaw experimentation emerges. Our analysis reveals specific parameter conditions under which a firm can consistently succeed in individual A/B tests while experiencing long-term performance decline across all dimensions.
To address this challenge, we propose a straightforward solution: implementing a positive hurdle rate for A/B test approval. This approach ensures that only innovations exceeding a specified threshold gain implementation, effectively balancing experimentation’s benefits against cumulative spillover costs. We derive the optimal hurdle rate at which the gain in the primary dimension equals the negative externality. We analyze how the optimal hurdle rate varies with the underlying parameters, providing practical guidance for experimental design.
@@
The literature addressing our specific focus is relatively small. Berman et al. (2018) examine how p-hacking in A/B testing can lead to false discoveries and impair experimental improvements. Azevedo et al. (2020) analyze how distributional characteristics, particularly fat tails, influence experimentation allocation strategies. McClellan (2022) explores how agency problems affect A/B test adoption mechanisms. While these studies provide valuable insights, they primarily focus on single-dimensional outcomes. Our work departs from this literature by explicitly modeling the interconnectedness of performance across multiple dimensions, revealing how improvements in one dimension can generate spillover effects in others.
@@ 2 Model Formulation
    $\{a_{t}:t\geq 1\}$ is a sequence of i.i.d. realizations of a random variable $A$, where $\mathbb{P}(A=u)=p_{u}$, $\mathbb{P}(A=v)=p_{v}$, and $p_{u}+p_{v}=1$.
@@
    $\{(u_{t},v_{t}):t\geq 1\}$ is a sequence of i.i.d. realizations of a random vector $(U,V)$, following a bivariate normal distribution with mean vector $m$ and covariance matrix $\Sigma$:
  3. (iii)
    Both $\mu_{u}$ and $\mu_{v}$ are negative.
  4. (iv)
@@
Under these conditions, the firm’s overall performance is a sum of i.i.d. random variables:
@@
Assumption 1 allow us to deliver the main message of the paper—the existence of seesaw experimentation and its mitigation via a positive hurdle rate—in the simplest and most standard framework. The conditions in this assumption can be relaxed without affecting the main points of the paper. In particular, in condition (i), $a_{t}$ can be correlated over time. For condition (ii), our results also apply to broader classes of distributions, including those whose values can on discrete values. Condition (iii) is not necessary in the sense that seesaw experimentation can still occur in scenarios where either $\mu_{u}$ or $\mu_{v}$ is non-negative while the other is negative, provided that $\mu_{u}+\mu_{v}<0$. But we assume this condition to reflect the fundamental business reality that random innovations typically do not improve secondary performance dimensions. Finally, condition (iv) is also not necessary in the sense that our main results hold even if the choice of innovation can depend on arbitrary realizations of outcomes in the past. Below, we characterize conditions under which seesaw experimentation occurs and analyze how to determine an optimal hurdle rate to address this issue.
3 Symmetric Case
@@
In the symmetric case, $\mathbb{E}[D(U+V)]<0$ if
$-1\leq\rho<2\left(\frac{-\mu}{\sigma}\right)M\left(\frac{-\mu}{\sigma}\right)-1,$
@@
where $M(\alpha)\coloneqq(1-\Phi(\alpha))/\phi(\alpha)$ is the Mills ratio of the standard normal distribution, with $\phi$ and $\Phi$ representing the probability density function and cumulative distribution function of the standard normal distribution, respectively.
Proposition 1 provides a sufficient condition for the phenomenon of seesaw experimentation. Since $\alpha M(\alpha)$ is an increasing function of $\alpha$ (see Lemma 1 in the Appendix), seesaw experimentation becomes more likely as the signal-to-noise ratio $\alpha=|\mu|/\sigma$ grows, whether through an increase in $|\mu|$ or a decrease in $\sigma$. The likelihood of seesaw experimentation also increases as the correlation $\rho$ becomes more negative. Moreover, as $(2\alpha M(\alpha)-1)$ monotonically increases from $-1$ to $1$ for $\alpha\in[0,\infty)$, seesaw experimentation can occur even with positive correlation.
Proposition 1 shows that even if A/B tests constantly improve the firm’s performance in the primary dimension, say in the department that fits with the firm’s strategic priority, the firm’s overall performance may decrease in the long run. This drop occurs when the negative spillover on the secondary dimension (other department) is sufficiently serious. In practice, there can be many solutions to this problem, including interdepartmental communication and coordination. But they often prove impractical given the high volume of experiments and the difficulty in measuring secondary dimension outcomes. Below, we propose a simple remedy that does not rely on interdepartmental coordination: implementing a positive hurdle rate $z$. Under this mechanism, an innovation in either dimension is adopted only if its direct effect in that dimension exceeds $z$, thereby creating a more stringent filter for innovation adoption.
Proposition 1 reveals a counterintuitive finding: while a firm may seemingly march from one victory to another through successful A/B tests, its overall performance may deteriorate over time. This decline stems from significant negative spillover effects on secondary dimensions that offset the gain in the primary dimension. Although traditional remedies such as interdepartmental coordination exist, they become unwieldy in practice due to the sheer volume of experiments and challenges in measuring the effects on the secondary dimension. We therefore propose a simpler solution: the implementation of a positive hurdle rate $z$. This approach requires that any innovation must demonstrate a benefit exceeding $z$ in the primary dimension before adoption, effectively raising the bar for implementation without requiring complex coordination across departments.
@@
In this case, the firm’s long-run average performance becomes:
@@
In the symmetric case, if $\rho\in(-1,1)$, then $f^{\prime}(0)>0$, $f(z)$ attains its maximum at $z^{*}$ and $f(z^{*})>0$, where
@@
Proposition 2 demonstrates that implementing a positive hurdle rate is strictly beneficial for the firm. The intuition is straightforward: When the firm sets a hurdle rate slightly above zero, it experiences two effects. First, there is a small loss in the primary dimension because some marginally beneficial innovations are not implemented. Since these are innovations that barely clear a zero hurdle rate, the foregone benefits are minimal. Second, there is a significant gain in the secondary dimension because the firm avoids implementing innovations that would have negative effects (in expectation) there. By setting a positive hurdle rate, the firm screens out potential innovations that might have small positive benefits in the primary dimension but large negative effects in the secondary dimension. This makes a positive hurdle rate optimal for the firm.
The proposition also determines the optimal hurdle rate and reveals two key relationships. First, the optimal hurdle rate is proportionally related to $\mu$ (the mean value in the secondary dimension). When $\mu$ is lower, the firm sets a higher hurdle rate. This makes intuitive sense as the firm wants to be more selective when secondary outcomes are likely to be less favorable.
@@
The optimal hurdle rate in Proposition 2 maximizes the firm’s long-term performance. A firm that also prioritizes short-term objectives would likely set a lower hurdle rate, accepting more innovations despite their potential negative secondary effects. This reflects the classic trade-off between short-term gains and long-term losses.
@@ 4 Extensions
In this section, we extend the results of Section 3 to asymmetric bivariate normal distributions and the multidimensional case.
4.1 Asymmetric Case
@@
In the asymmetric case, $g(0,0)<0$ if $-1\leq\rho<\min(\rho_{1},\rho_{2})$, where
@@ 4.2 Multidimensional Case
Assume $(X_{1},\ldots,X_{n})$ follows a multivariate normal distribution with $\mathbb{E}[X_{i}]=\mu<0$ and $\mathrm{Var}[X_{i}]=\sigma^{2}$ for all $i$, and $\mathrm{Cov}[X_{i},X_{j}]=\rho\sigma^{2}$ for all $i\neq j$. Then, $h(z)<0$ if
@@
Under the multivariate normal assumption of Proposition 5, if $-1/(n-1)<\rho<1$, then $h^{\prime}(0)>0$, $h(z)$ attains its maximum at $z^{*}$, and $h(z^{*})>0$, where
$z^{*}=\frac{(n-1)(\rho-1)}{1+(n-1)\rho}\mu.$

### v3

@@
Seesaw Experimentation: A/B Tests with Spillovers thanks: Contact information: Jin Li (jli1@hku.hk), Ye Luo (kurtluo@hku.hk), Faculty of Business and Economics, The University of Hong Kong, Pokfulam Road, Hong Kong SAR. Xiaowei Zhang (xiaoweiz@ust.hk), Department of Industrial Engineering and Decision Analytics, The Hong Kong University of Science and Technology, Clear Water Bay, Hong Kong SAR. We thank Shan Huang and Ruohan Zhan for helpful discussions.
@@
(December 31, 2024)
This paper investigates how a firm’s performance can decline even when consistently implementing innovations validated through A/B testing. We introduce the concept of seesaw experimentation, where successful innovations enhance performance in measured primary metrics but generate negative spillover effects in unmeasured secondary dimensions, ultimately reducing overall performance. We identify the conditions under which seesaw experimentation occurs. We also propose a simple solution to address it: implementing a positive hurdle rate for A/B tests. We derive the optimal hurdle rate for both the normal distribution and the fat-tailed distributions.
@@ 1 Introduction
A/B testing has become an important tool for modern business, enabling firms to drive innovation, improve user experiences, and increase revenue growth (Kohavi et al., 2020, Koning et al., 2022). Major tech firms such as Amazon, Google, Meta, and Microsoft deploy thousands of such tests annually, embedding data-driven decision making into their innovation processes (Thomke, 2020, Azevedo et al., 2020).
While existing research typically examines A/B tests in isolation, these experiments frequently generate significant spillover effects [^1] We use spillover, rather than negative spillover, for expositional ease. that ripple across an organization’s ecosystem of products, channels, and performance metrics. These effects are particularly pronounced on large digital platforms. At the WeChat platform, for instance, enhancing Video Accounts can substantially decrease engagement with Official Accounts as users redistribute their limited attention. The spillover effects extends beyond digital interactions: improvements that drive online sales growth often cannibalize offline revenue, and strategies that maximize advertising revenue frequently degrade user experience. These examples underscore a fundamental challenge: improvements in one area of business operations can trigger unintended consequences in seemingly unrelated domains. As externalities are prevalent within organizations, so are the spillover effects.
These spillover effects create what we term seesaw experimentation, where the overall performance of a firm deteriorates despite continuous improvements through innovations in current strategic priorities. The name reflects its oscillating nature. Like a seesaw, advances in one dimension often diminish performance in others. Firms implement innovations that boost their primary metrics—whether in specific products, channels, or KPIs. Yet these same innovations generate hidden or underweighted losses in secondary dimensions. Because business environments and strategic priorities constantly evolve, yesterday’s focus on gross merchandise value may shift to user satisfaction tomorrow. Each shift triggers new waves of innovations that may inadvertently reverse previous gains. The pattern creates a dangerous illusion of progress. Success metrics in current focal areas shine brightly. Meanwhile, the firm’s overall performance quietly degrades.
Managing spillover effects in A/B testing poses a formidable challenge. This is especially true for large organizations running numerous simultaneous experiments. Some firms try to address single-product spillovers through managerial coordination. For instance, user growth teams may challenge revenue-boosting innovations that hurt customer satisfaction. But such efforts often fall short. Several factors drive this coordination failure. First, many potential effects—especially long-term impacts—defy easy quantification. Second, the sheer volume of concurrent experiments makes comprehensive coordination impractical. Third, cross-product and cross-channel effects often remain hidden from view. Despite widespread recognition of their importance for platform health, spillover effects rarely factor into experimental design [^2] Communications with experts at a leading Chinese platform. This creates a persistent gap between the ideal of holistic testing and the reality of fragmented experimentation.
The purpose of this paper is to study A/B testing in the presence of spillover effects, with particular attention to the emergence and implications of seesaw experimentation. We investigate the conditions under which firms, despite achieving continuous improvements in primary performance metrics, may experience deteriorating overall outcomes. Our analysis focuses on two key questions: First, under what conditions does seesaw experimentation emerge as firms improve their primary strategic objectives? Second, how can organizations effectively manage these spillover effects without relying on complex cross-managerial coordination?
We develop a bivariate normal distribution framework to demonstrate the emergence of seesaw experimentation. Our model identifies conditions where firms achieve consistent A/B test successes yet face declining overall performance. Two key factors drive this phenomenon. First, a higher signal-to-noise ratio (the absolute value of mean over standard deviation) increases the likelihood of seesaw experimentation. Second, this effect strengthens as the correlation between dimensions becomes more negative.
Our analysis yields a simple yet powerful solution to manage spillover: implement a positive hurdle rate for A/B test approval. This approach only implements innovations that exceed a specified performance threshold. We derive the optimal hurdle rate where the positive gains in the primary dimension equal the negative externalities in the secondary dimension. Our framework provides practical guidance by showing how this optimal rate varies with key distribution parameters. This solution effectively balances innovation benefits against spillover costs without requiring complex coordination mechanisms.
@@
The literature addressing our specific focus is relatively small. Berman et al., (2018) examine how p-hacking in A/B testing can lead to false discoveries and impair experimental improvements. Azevedo et al., (2020) analyze how distributional characteristics, particularly fat tails, influence experimentation allocation strategies. [^3] We extend our analysis to fat-tailed distributions in Section 4. McClellan, (2022) explores how agency problems affect A/B test adoption mechanisms. While these studies provide valuable insights, they primarily focus on single-dimensional outcomes. Our work departs from this literature by explicitly modeling the interconnectedness of performance across multiple dimensions, revealing how improvements in one dimension can generate spillover effects in others.
@@ 2 Model Formulation
    $\{a_{t}:t\geq 1\}$ is a sequence of i.i.d. realizations of a random variable $A$, where $\mathbb{P}(A=u)=p_{u}$, $\mathbb{P}(A=v)=p_{v}$, and $p_{u}+p_{v}=1$. $\{(u_{t},v_{t}):t\geq 1\}$ is a sequence of i.i.d. realizations of a random vector $(U,V)$.
@@
Under Assumption 1, the firm’s overall performance is a sum of i.i.d. random variables:
@@
For simplicity, we start by considering bivariate normal distributions. We extend our results to a wider class of distributions later.
Assumption 2.
  1. (i)
    $(U,V)$ follows a bivariate normal distribution with mean vector $m$ and covariance matrix $\Sigma$:
  2. (ii)
    Both $\mu_{u}$ and $\mu_{v}$ are negative.
Assumptions 1 and 2 allow us to deliver the main message of the paper—the existence of seesaw experimentation and its mitigation via a positive hurdle rate—in the simplest and most standard framework. The conditions in these assumptions can be relaxed without affecting the main points of the paper. In particular, in Assumption 1(i), $a_{t}$ can be correlated over time, and Assumption 1(ii) is also not necessary in the sense that our main results hold even if the choice of innovation can depend on arbitrary realizations of outcomes in the past. For Assumption 2(i), our results also apply to broader classes of distributions, including those having fat tails and those whose values can on discrete values. Assumption 2(ii) is not necessary in the sense that seesaw experimentation can still occur in scenarios where either $\mu_{u}$ or $\mu_{v}$ is non-negative while the other is negative, provided that $\mu_{u}+\mu_{v}<0$. But we assume this condition to reflect the fundamental business reality that random innovations typically do not improve secondary performance dimensions. Below, we characterize conditions under which seesaw experimentation occurs and analyze how to determine an optimal hurdle rate to address this issue.
3 Base Model
@@
Suppose Assumptions 1 and 2 hold and $(U,V)$ follows a symmetric distribution. If
$-1\leq\rho<2\left(\frac{-\mu}{\sigma}\right)M\left(\frac{-\mu}{\sigma}\right)-1,$ (1)
@@
then $\mathbb{E}[D(U+V)]<0$, where $M(\alpha)\coloneqq(1-\Phi(\alpha))/\phi(\alpha)$ is the Mills ratio of the standard normal distribution, with $\phi$ and $\Phi$ representing the probability density function and cumulative distribution function of the standard normal distribution, respectively.
Proposition 1 shows that seesaw experimentation can occur. That is, while a firm may seemingly improve its performance constantly through successful A/B tests, its overall performance may deteriorate over time. This decline stems from significant negative spillover effects on secondary dimensions that offset the gain in the primary dimension.
Proposition 1 also provides a sufficient condition for seesaw experimentation to occur. One factor is the signal-to-noise ratio $\alpha=|\mu|/\sigma$. We show in the Appendix (Lemma 1) that $\alpha M(\alpha)$ is an increasing function of $\alpha$, implying that the sufficient condition is more likely to be satisfied when signal-to-noise ratio increases. This can happen in two ways: either through an increase in $|\mu|$ or a decrease in $\sigma$. The intuition behind this is straightforward. When $|\mu|$ is larger, the negative impact on the secondary dimension becomes more pronounced. Alternatively, when $\sigma$ is smaller, there’s less variability in outcomes, which typically results in smaller average benefits in the primary dimension. Both scenarios make seesaw experimentation more likely to occur.
Another key factor is the correlation coefficient $\rho$ between the two dimensions. As $\rho$ becomes more negative, seesaw experimentation becomes more likely, reflecting situations where improvements in the primary dimension are increasingly linked to negative effects in the secondary dimension. While one might intuitively expect seesaw experimentation to occur only with negative correlation, our analysis reveals that this is not true. Because $(2\alpha M(\alpha)-1)$ monotonically increases from $-1$ to $1$ for $\alpha\in[0,\infty)$, our sufficient condition indicates that seesaw experimentation can emerge even with positive correlation. This counterintuitive result occurs, for example, when $|\mu|$ is large, meaning the average negative effect on the secondary dimension is sufficiently strong to override the positive correlation between dimensions.
How to deal with seesaw experimentation? The traditional approach relies on inter-departmental coordination. For instance, customer satisfaction executives might veto an A/B test that boosts revenue but diminishes customer experience. However, this approach faces several challenges. The relative importance of different dimensions often involves subjective judgment, and outcomes can vary depending on individual executives’ perspectives and personalities. Moreover, when companies carry out thousands of A/B tests, the sheer volume of experiments and the difficulty in measuring secondary dimension effects make coordination costs prohibitively high.
Below, we propose a coordination-free solution: the implementation of a positive hurdle rate $z$. This approach requires that any innovation must demonstrate a benefit exceeding $z$ in the primary dimension before adoption, effectively raising the bar for implementation without requiring complex coordination across departments.
@@
When a positive hurdle $z$ is implemented, the firm’s long-run average performance becomes:
@@
Suppose Assumptions 1 and 2 hold and $(U,V)$ follows a symmetric distribution. If $\rho\in(-1,1)$, then $f^{\prime}(0)>0$, $f(z)$ attains its maximum at $z^{*}$ and $f(z^{*})>0$, where
@@
Proposition 2 demonstrates that implementing a positive hurdle rate is strictly beneficial to the firm. The intuition is straightforward: When the firm sets a hurdle rate slightly above zero, there are two conflicting effects. First, there is a small loss in the primary dimension because some marginally beneficial innovations are not implemented. Note that the foregone benefits are small for these innovations. Second, there is a significant gain in the secondary dimension because the firm avoids implementing innovations that would have negative effects (in expectation) there. By setting a positive hurdle rate, the firm screens out potential innovations that might have small positive benefits in the primary dimension but large negative effects in the secondary dimension. This makes a positive hurdle rate beneficial.
This proposition also determines the optimal hurdle rate and reveals two key relationships. First, the optimal hurdle rate is proportionally related to $\mu$ (the mean value in the secondary dimension). When $\mu$ is lower, the firm sets a higher hurdle rate. This makes intuitive sense, as the firm wants to be more selective when secondary outcomes are likely to be less favorable.
@@
Finally, note that the optimal hurdle rate in Proposition 2 maximizes the firm’s long-term performance. A firm that also prioritizes short-term objectives should set a lower hurdle rate, accepting more innovations despite their potential negative secondary effects. This reflects the classic trade-off between short-term gains and long-term losses.
@@ 4 Extensions
This section extends the results of Section 3 in three directions: asymmetry, multidimensionality, and fat tails. For clarity of managerial insights, we examine each extension separately rather than analyzing a fully general model combining all features.
4.1 Asymmetry
@@
Suppose Assumptions 1 and 2 hold. If $-1\leq\rho<\min(\rho_{1},\rho_{2})$, then $g(0,0)<0$, where
@@ 4.2 Multidimensionality
Assumption 3.
  1. (i)
    $(X_{1},\ldots,X_{n})$ follows a multivariate normal distribution with mean vector $m$ and covariance matrix $\Sigma$:
    $m=\begin{pmatrix}\mu\\ \mu\\ \vdots\\ \mu\end{pmatrix}\quad\mbox{and}\quad\Sigma=\begin{pmatrix}\sigma^{2}&\rho\sigma^{2}&\cdots&\rho\sigma^{2}\\ \rho\sigma^{2}&\sigma^{2}&\cdots&\rho\sigma^{2}\\ \vdots&\vdots&\ddots&\vdots\\ \rho\sigma^{2}&\rho\sigma^{2}&\cdots&\sigma^{2}\end{pmatrix}.$
  2. (ii)
    $\mu<0$.
@@
Suppose Assumptions 1 and 3 hold. Then, $h(0)<0$ if
@@
Suppose Assumptions 1 and 3 hold. If $-1/(n-1)<\rho<1$, then $h^{\prime}(0)>0$, $h(z)$ attains its maximum at $z^{*}$, and $h(z^{*})>0$, where
$z^{*}=\frac{(n-1)(\rho-1)}{(n-1)\rho+1}\mu.$
@@
4.3 Fat Tails
Azevedo et al., (2020) found that A/B test results frequently display fat tails that cannot be adequately modeled by normal distributions. Building on this insight, we extend our analysis to bivariate $t$ distributions. This class of distributions extends bivariate normal distributions by adding a parameter that controls tail thickness. Unlike normal distributions, their tail probabilities decay polynomially, with the added parameter determining the decay rate. [^6] While Pareto distributions also model heavy tails effectively, their bivariate extensions have limited flexibility: they are not defined over the entire real plane, and their correlation structure is constrained by the tail thickness parameter rather than being independently specified (Arnold, 2015).
Assumption 4.
  1. (i)
    $(U,V)$ follows a bivariate $t$ distribution with mean vector $m$, scale matrix $\Sigma$, and degrees of freedom $\delta$:
    $m=\begin{pmatrix}\mu\\ \mu\end{pmatrix}\quad\mbox{and}\quad\Sigma=\begin{pmatrix}\sigma^{2}&\rho\sigma^{2}\\ \rho\sigma^{2}&\sigma^{2}\end{pmatrix}.$
  2. (ii)
    $\mu<0$ and $\delta>2$.
The parameter $\delta$ controls the tail thickness of the distribution, with smaller values of $\delta$ corresponding to heavier tails. When $\delta=2$, the distribution lacks finite second moments, making covariance and correlation undefined. For $\delta>2$, the scale matrix $\Sigma$ equals the covariance matrix of $(U,V)$, where $\sigma^{2}$ represents the marginal variance and $\rho$ the correlation. As $\delta\to\infty$, the distribution converges to a bivariate normal with the same mean vector and covariance matrix, eliminating the fat tails.
Proposition 7.
Suppose Assumptions 1 and 4 hold. If
$-1\leq\rho<2\left(\frac{-\mu}{\sigma}\right)W\left(\frac{-\mu}{\sigma}\right)-1,$ (5)
then $f(0)<0$, where
$W(\alpha)=\left(\frac{\delta-2}{\delta}\right)\left(\frac{1-T_{\delta}(\alpha;0,1)}{t_{\delta-2}\left(\alpha;0,\frac{\delta}{\delta-2}\right)}\right),$
$t_{\delta}(\alpha;\mu,\sigma^{2})$ and $T_{\delta}(\alpha;\mu,\sigma^{2})$ denote the probability density function and cumulative distribution function, respectively, of a univariate $t$ distribution with mean $\mu$, variance $\sigma^{2}$, and degrees of freedom $\delta$.
In equation (5), the maximum correlation permitting seesaw experimentation is $12\alpha W(\alpha)-$ 1, where $\alpha=-\mu/\sigma$. Figure 2 illustrates this threshold as a function of $\alpha$ for different values of $\delta$. As $\delta$ increases, this threshold increases for any given $\alpha$. This indicates that, holding marginal mean ($\mu$) and variance ($\sigma^{2}$) constant, lighter tails expand the range of correlations that permit seesaw experimentation, making such phenomenon more likely to occur. This result has an intuitive explanation: with lighter tails, experimentation offers more limited upside potential. Consequently, the negative spillover effects in the secondary dimension are more likely to outweigh the positive gains in the primary dimension.
Figure 2: Maximum Correlation Leading to Seesaw Experimentation
Remark 1.
The standard $t$ distribution with $\delta$ degrees of freedom converges to the standard normal distribution as $\delta\to\infty$. Thus,
$W(\alpha)=\left(\frac{\delta-2}{\delta}\right)\left(\frac{1-T_{\delta}(\alpha;0,1)}{t_{\delta-2}\left(\alpha;0,\frac{\delta}{\delta-2}\right)}\right)\to\frac{1-T_{\infty}(\alpha;0,1)}{t_{\infty}(\alpha;0,1)}=\frac{1-\Phi(\alpha)}{\phi(\alpha)}=M(\alpha).$
Hence, the upper bound in equation (5) for the bivariate $t$ distribution converges to the upper bound in (1) for the base model (bivariate normal distribution).
Proposition 8.
Suppose Assumptions 1 and 4 hold. If $\rho\in(-1,1)$, then $f^{\prime}(0)>0$, $f(z)$ attains its maximum at $z^{*}$ and $f(z^{*})>0$, where
$z^{*}=\frac{\rho-1}{\rho+1}\mu.$
Note that the choice of the optimal hurdle rate is the same as that in the normal distribution. The underlying intuition, illustrated in Figure 1, holds true: the optimal hurdle rate equates the expected negative externality in the secondary dimension with the gain in the primary dimension. Both bivariate normal and bivariate $t$ distributions share the same expression for the expected negative externality: $\mathbb{E}[-V|U=u]=-\rho\mu-(1-\rho)\mu$.